a car leaves LA traveling north at 59 km/ph. another car leaves LA an hour later on the same road traveling 84 km/ph. how far from LA will the second car overtake the first car?
Answer by stanbon(75887)

a car leaves LA traveling north at 59 km/ph. another car leaves LA an hour later on the same road traveling 84 km/ph. how far from LA will the second car overtake the first car?
------
Draw the picture.
Notice that distance for the 2 cars is the same when one overtakes the other.
---
1st Car DATA:
rate = 59 kmh ; time = x hrs ; distance = 59x km
-----
2nd Car DATA:
rate = 84 kmh ; time = x-1 hrs ; distance = 84(x-1) km
----
Equation:
distance = distance
59x = 84x - 84
25x = 84
x = 3.36 hrs
x-1 = 2.36 hrs
Distance of 2nd car: 2.36*84 = 198.24 km
==============================================
Cheers,
Stan H.